40x4gx403=bw3E

60x4gx603=bw3E

Find E and g simultaneously




1
Expert's answer
2021-08-12T06:19:29-0400

There is no solution:


"\frac{40\cdot4g\cdot40^3}{60\cdot4g\cdot60^3}=\frac{32}{81}\\
\frac{bw^3E}{bw^3E}=1"